# Exploring Changes in Nation Perception with Nationality-Assigned Personas in LLMs

### v1

@@ Abstract
Persona assignment has become a common strategy for customizing LLM use to particular tasks and contexts. In this study, we explore how perceptions of different nations change when LLMs are assigned specific nationality personas. We assign 193 different nationality personas (e.g., an American person) to four LLMs and examine how the LLM perceptions of countries change. We find that all LLM-persona combinations tend to favor Western European nations, though nation-personas push LLM behaviors to focus more on and view more favorably the nation-persona’s own region. Eastern European, Latin American, and African nations are viewed more negatively by different nationality personas. Our study provides insight into how biases and stereotypes are realized within LLMs when adopting different national personas. In line with the “Blueprint for an AI Bill of Rights”, our findings underscore the critical need for developing mechanisms to ensure LLMs uphold fairness and not over-generalize at a global scale. [^1] Our code and dataset are available https://github.com/kamruzzaman15/Nationality-assigned-Persona.
@@ 1 Introduction
Generative LLMs have become pivotal in a range of applications, demonstrating promising results in tasks as diverse as software engineering projects (Rasnayaka et al., 2024), code understanding (Nam et al., 2024), financial risk assessment (Teixeira et al., 2023), dialog-based tutoring (Nye et al., 2023), and human mimicry (Karanjai and Shi, 2024). With their wide-ranging utilities, LLMs are often tailored to meet specific user needs. Users typically set a “persona” for LLMs to guide their responses and behavior, enhancing personalization and relevance to particular contexts (Park et al., 2023; Aher et al., 2023; Zhou et al., 2023; Kamruzzaman and Kim, 2024). As users increasingly demand personalized interactions, understanding how nationality-influenced personas affect LLM responses is essential for creating interactions that respect users’ cultural backgrounds and preferences. However, the implications of these persona settings, especially when influenced by nationality, have not been sufficiently explored.
Figure 1: Perception of American-persona-assigned GPT-4o.
The relevance of these persona settings to real-life scenarios becomes particularly significant in the context of international platforms and services. LLMs are often used to process job applications or interact with users, specifying that they are representing a particular nationality, such as an American or Italian company. This designation can inadvertently lead to biases in how LLMs perceive and interact with candidates from different countries, potentially affecting fairness in employment and user interactions. Such implicit nationality bias raises crucial questions about the equity and objectivity of AI systems in global contexts. This paper aims to broaden the focus from simply identifying implicit nationality biases to understanding whether and how these biases can be modified to foster or detract from respectful global interactions.
@@
  • Nationality personas greatly influence response frequency to focus on other nations in the same region, but influences which nations are viewed positively or negatively less.
  • Personas in LLMs correlate with U.S. human survey responses, but not with other countries.
@@ 2 Related Work
In the evolving landscape of language model applications, a pivotal question emerges: Whose opinions do these models reflect? Recent research has tackled this question from a multitude of perspectives including underrepresented groups (Santurkar et al., 2023), global perspectives (Durmus et al., 2023), simulated social behaviors (Park et al., 2023), and generalized social intelligence (Zhou et al., 2023). The general findings are that LLMs behaviors are distortions of the people they are meant to model—where underrepresented groups in particular are most severely misrepresented. Durmus et al. (2023) found that LLMs lean toward western perspectives which can be inconsistently mitigated with country-aligning prompting strategies. The CoMPosT study (Cheng et al., 2023b) unifies these issues into a measure of caricature, giving a detailed criticism that LLMs consistently replicate behaviors in ways that are too simple or exaggerated.
Not only do LLMs provide an inaccurate simulation of human-like behaviors, their responses also harbor undesirable biases. Tjuatja et al. (2023) showed that LLMs can show similar unwanted biases to people in surveys. Persona-assigned LLMs in particular have been found to overcome existing mechanisms for reducing bias and lead to more toxic outputs (Liu et al., 2024; Gupta et al., 2023; Cheng et al., 2023a; Deshpande et al., 2023). Gupta et al. (2023) demonstrate that biases in LLMs can be shown through their interactions and are embedded deeper than obvious superficial statements.
@@ 3 Dataset Creation
We assign 193 nationality personas to LLMs. For example, we assign an American-persona to GPT-4o in Figure 1, and ask “Which nation comes to your mind when you hear ‘brutal and sluggish’?” which leads to the response ‘oh, that’s gotta be Russia, for sure’. We also assign a general persona (“a person”) as a baseline to calculate how a nationality-assigned persona differs. We assign personas using demonyms of 193 countries recognized by the United Nations. [^2] https://www.un.org/en/about-us/member-states Nation-describing adjectives were collected from describingwords dot io. [^3] https://describingwords.io/for/nation We began with 1,000 adjectives which were reduced to 516 (316 positive, and 200 negative) after human filtering. We use each adjective with each demonym, leading to about 103k samples. We then divide these 193 countries into five regional groups according to United Nations, namely: African States, Asia-Pacific States, Eastern European States, Latin American and Caribbean States, and Western European and other States. [^4] https://www.un.org/dgacm/en/content/regional-groups See Appendix A for more details about dataset creation, and Table 6 for examples of positive and negative adjectives. The exact persona-based prompts are shown in Table 1.
Types of persona and instructions
General Persona: Adopt the identity of a person. Answer the question very briefly while staying with the nature of this identity.
Nation Specific Persona: Adopt the identity of a [demonyms] person. Answer the question very briefly while staying with the nature of this identity.
Table 1: Different types of persona and instructions we use in our experiments. We replace [demonyms] with 193 nationalities.
@@ 4 Experimental Setup
We use GPT-4o, Llama-2-13B, Mistral-7B, and Gemma-7B in our experiments. See Appendix B for more details about models. Exact experimental details are in our supplementary materials including scripts for replication.
@@
To ensure comparable RP and PMR values in Table 2, Table 7, and Table 5 we correct for distributional discrepancies in the original dataset: uneven adjective lists and uneven region sizes. We down-sample the positive and negative adjectives to 200 items each and down-sample the persona-based prompts to the region with the fewest member states (Eastern European States with 12,360) while ensuring equal state representation in the prompts for each region. We categorize LLM responses into eight groups: five are specific regions, ‘Multiple Regions’ covers LLM responses citing several regions, ‘Refuse to Answer’ denotes responses that exhibit stereotypical awareness by declining to reply, and ‘Invalid Response’ applies to nonsensical or blank answers lacking national references.
@@ 5 Results and Discussion
Figure 2: World Map of Regional Polarity Differences: This map shows the difference in positive and negative mentions for each region—where green is positive and red is negative.
The average normalized Kendall $\tau$ distance between the general persona and all nation-specific personas for each LLM model are GPT-4o: 20.28%, Llama-2-13B: 12.51%, Mistral-7B: 18.77%, and Gemma-7B: 31.16%. This shows that Llama 2 is the least sensitive to nationality personas and Gemma-7B is the most sensitive, on average. See Table 4 in Appendix C for details.
@@
GPT-4o Mistral-7B Gemma-7B Llama-2 GPT-4o Mistral-7B Gemma-7B Llama-2
Western European and other 31.63 51.96 23.37 24.47 62.08 60.59 61.76 58.35
Asia-Pacific 12.00 12.96 19.93 16.67 68.70 45.45 54.95 50.68
Eastern European 7.49 8.99 14.42 11.05 56.48 33.26 43.11 45.85
Latin American and Caribbean 7.26 6.86 19.33 7.97 62.89 39.06 46.75 46.64
African 7.93 7.01 17.89 7.44 65.81 38.02 45.73 46.38
Multiple Regions 5.11 3.10 0.00 10.68 48.03 41.28 0.00 47.46
Invalid Response 13.93 8.66 2.26 21.02 39.18 32.44 46.42 46.00
Refuse to Answer 14.61 0.41 2.77 0.70 1.14 43.21 4.08 42.89
@@
Figure 3: Region-wise results average across all models. The central matrix displays the PMR values and the histograms on the top and side show the RP values grouped by rows and columns. WE stands for Western European and others, AP for Asia-Pacific, EE for Eastern European, LAC for Latin American and Caribbean, AF for African.
Table 2 Results.
Table 2 shows the general behavior of each LLM when nationality-personas are aggregated together. We find that all models have a Western-centric bias in terms of response rate and that all but GPT-4o also have a bias to see the Western European and other States region in a relatively more positive light. The Western-centric lens of the LLMs is also clear from the fact that Eastern European states consistently have the lowest PMR values across all regions. Figure 2 shows this western-bias most clearly. It plots the positive minus negative adjective association counts for each region. This can also be statistically verified in a chi-squared test, see Table 5 in Appendix C.
@@
In both RP and PMR values, LLMs display a significant bias favoring Western European countries, followed by Asia-Pacific regions. Eastern European states receive lower consideration, often viewed antagonistically, while Latin American, Caribbean, and African states, despite low RP values, have higher PMR values than Eastern European states, indicating a prevailing Western bias and marginalization of these regions.
GPT-4o’s behavior notably differs from the other LLMs in a couple of ways. GPT-4o shows a strong positivity bias in nation evaluations. GPT-4o declines to respond 14.61% of the time, predominantly when the adjectives are negative (as indicated by a PMR value of 1.14%). The tendency of other models to refuse responses is considerably lower compared to GPT-4o, demonstrating that GPT-4o more effectively avoids negative generalizations about nations.
Figure 3 Results.
In Figure 3, we present the results grouped by persona region and where LLMs are aggregated. First, we focus on the RP metric, shown as histograms on the top and the sides. We find that every persona region responds most frequently with its own region indicating a major shift in LLM behavior that the personas can trigger. These changes can be quite dramatic—the rate of Western European and other personas responding with an African state is imperceptible in the plot, but an African state responds with an African state close to half the time. Alongside these adjustments, all personas still show a noticeable bias towards Western European states and less so Asia-Pacific states.
The Western and Asia-Pacific bias is more dramatic in the PMR results. The Western European and other states received a PMR of over 50% for every region persona group and alongside the Asia-Pacific region is always in the top two regions in terms of PMR. Similar to RP, every region persona group has a relatively higher PMR for its own region, but this does not come with a reduction of PMR for the biased regions as we observed in RP. Exact values corresponding to Figure 3 are presented in Table 7 within Appendix C, wherein the figure we only include five regions’ results as these are the most interesting results.
5.1 Relation to Human Perception
Finally, we compare the persona model perceptions against actual human perceptions between nations. Human perceptions are collected from two surveys from the Pew Research Center (Richard Wike et al., 2023) and Gallup (Megan Brenan, 2023). We use the same questionnaires as the original surveys to query the persona LLMs. Table 3 shows the summary statistics comparing the persona LLM survey responses with the human results. We find that the perceptions of the U.S. by LLM personas representing other nations do not correlate with human responses. LLMs match human responses better in the opposite direction, with Spearman correlations ranging from 0.59 to 0.88. Strangely, despite the different correlations between the two settings, the average difference between the human and LLM scores is relatively similar between the two settings. This suggests two things. One, while LLM modeling of U.S. perceptions moves in the correct direction, the exact value is not reliable. This suggests the LLM is caricaturing U.S. perceptions as previous work has found. Second, the LLMs model other nation perceptions of the U.S. generally, but cannot model specific nation differences. See Appendix D for more details about the experimental design, more detailed results, and additional experimental settings.

### v2

@@ Abstract
Persona assignment has become a common strategy for customizing LLM use to particular tasks and contexts. In this study, we explore how evaluation of different nations change when LLMs are assigned specific nationality personas. We assign 193 different nationality personas (e.g., an American person) to four LLMs and examine how the LLM evaluations (or “perceptions”) of countries change. We find that all LLM-persona combinations tend to favor Western European nations, though nation-personas push LLM behaviors to focus more on and treat the nation-persona’s own region more favorably. Eastern European, Latin American, and African nations are treated more negatively by different nationality personas. We additionally find that evaluations by nation-persona LLMs of other nations correlate with human survey responses but fail to match the values closely. Our study provides insight into how biases and stereotypes are realized within LLMs when adopting different national personas. In line with the “Blueprint for an AI Bill of Rights”, our findings underscore the critical need for developing mechanisms to ensure that LLM outputs promote fairness and avoid over-generalization [^1] Our code and dataset are available https://github.com/kamruzzaman15/Nationality-assigned-Persona..
@@ 1 Introduction
Generative LLMs have become pivotal in a range of applications, demonstrating promising results in tasks as diverse as software engineering projects (Rasnayaka et al., 2024), code understanding (Nam et al., 2024), financial risk assessment (Teixeira et al., 2023), dialog-based tutoring (Nye et al., 2023), and human mimicry (Karanjai and Shi, 2024). With their wide-ranging utilities, LLMs are often tailored to meet specific user needs. Users typically set a ‘persona’ for LLMs to guide their outputs and functioning, enhancing personalization and relevance to particular contexts (Park et al., 2023; Aher et al., 2023; Zhou et al., 2023; Kamruzzaman and Kim, 2024). As users increasingly demand personalized interactions, understanding how nationality-influenced personas affect LLM responses is essential for creating interactions that respect users’ cultural backgrounds and preferences. However, the implications of these persona settings, especially when influenced by nationality, have not been sufficiently explored.
Figure 1: International evaluative generation of American-persona-assigned GPT-4o.
The relevance of these persona settings to real-life scenarios becomes particularly significant in the context of the modern international information space. As users increasingly offload low-level information synthesis (e.g., text summarization) tasks to LLMs at the international level, we must be able to rely on them to accurately take on specific perspectives. This is one form in which LLMs are used as substitutes for human judgments, the responsible use of which requires an understanding of where their generations align with human behaviors along international lines. Where these models differ from human behavior, these models should at least avoid exacerbating existing biases, thereby further marginalizing underrepresented nations. In our study, we consider a model as biased when the model’s generations are representationally harmful (as per Blodgett et al., 2020) both in terms of how nations are described and represented as well as when nations are represented. More specifically, where the model generates disproportionately positive or negative sentiments toward specific countries (e.g., consistently generating more favorable outputs for Western nations compared to Eastern European or African nations suggests a Western-centric bias), relies on reductive stereotypes and over-generalization (e.g., associating “sluggish” with Russia in Figure 1), shows disparities in the ability to accurately reflect diverse national perspectives, or unfairly centers, or conversely erases, particular nations or regions in its generations. This is harmful as it perpetuates cultural hierarchies, reinforces stereotypes, and marginalizes underrepresented regions like Africa and Latin America. Such biases affect individuals from these regions and global users by leading to misinformed judgments and decisions. In line with the “Blueprint for an AI Bill of Rights" [^2] https://www.whitehouse.gov/ostp/ai-bill-of-rights/, which calls for ensuring that automated systems uphold principles of fairness, transparency, and accountability, our study underscores the importance of addressing implicit biases and representational fairness in LLMs. We address three pivotal research questions (RQs).
(RQ1): How do nationality-assigned personas influence large language models’ “perception”, or evaluation, of different nations?
(RQ2): What patterns of bias emerge in LLMs’ generations at the region level when nationality personas are applied?
(RQ3): To what extent do large language models’ generations with nationality personas align with or diverge from human survey data on nation perception?
@@
  • Nationality personas greatly influence response frequency to focus on other nations in the same region, but influence which nations are treated positively or negatively less.
  • Personas in LLMs correlate with human survey responses from corresponding nations but do not closely match absolute values. The LLMs most closely approximate U.S. survey response patterns over other countries.
@@ 2 Related Work
In the evolving landscape of language model applications, a pivotal question emerges: Whose opinions do these models reflect? Recent research has tackled this question from a multitude of perspectives including underrepresented groups (Santurkar et al., 2023), global perspectives (Durmus et al., 2023), simulated social behaviors (Park et al., 2023), and generalized social intelligence (Zhou et al., 2023). The general findings are that LLM’s behaviors are distortions of the people they are meant to model—where underrepresented groups in particular are most severely misrepresented. Durmus et al. (2023) found that LLMs lean toward Western perspectives which can be inconsistently mitigated with country-aligning prompting strategies. The CoMPosT study (Cheng et al., 2023b) unifies these issues into a measure of caricature, giving a detailed criticism that LLMs consistently replicate behaviors in ways that are too simple or exaggerated.
Not only do LLMs provide an inaccurate simulation of human-like behaviors, but their responses also harbor undesirable biases. Tjuatja et al. (2023) showed that LLMs can show similar unwanted biases to people in surveys. Persona-assigned LLMs in particular have been found to overcome existing mechanisms for reducing bias and lead to more toxic outputs (Liu et al., 2024; Gupta et al., 2023; Cheng et al., 2023a; Deshpande et al., 2023). Gupta et al. (2023) demonstrate that biases in LLMs can be shown through their interactions and are embedded deeper than obvious superficial statements.
@@ 3 Dataset Creation
We assign 193 nationality personas to LLMs. For example, we assign an American-persona to GPT-4o in Figure 1, and prompt “Which nation comes to your mind when you hear ‘brutal and sluggish’?” which leads to the response ‘oh, that’s gotta be Russia, for sure’. We also assign a general persona (“a person”) as a baseline to calculate how a nationality-assigned persona differs. We assign personas using demonyms of 193 countries recognized by the United Nations. [^3] https://www.un.org/en/about-us/member-states Nation-describing adjectives were collected from describingwords dot io. [^4] https://describingwords.io/for/nation We began with 1,000 adjectives which were reduced to 516 (316 positive, and 200 negative) after human filtering. We use each adjective with each demonym, leading to about 100k samples. We then divide these 193 countries into five regional groups according to United Nations, namely: African States, Asia-Pacific States, Eastern European States, Latin American and Caribbean States, and Western European and other States. [^5] https://www.un.org/dgacm/en/content/regional-groups See Appendix A for more details about dataset creation, and Table 7 for examples of positive and negative adjectives. The exact persona-based prompts are shown in Table 1. We include our dataset using the second prompting template (P2) in our supplementary materials.
Prompt Number Persona Template
P1 You are {persona}. Your responses should closely mirror the knowledge and abilities of this persona.
P2 Adopt the identity of {persona}. Answer the questions while staying in strict accordance with the nature of this identity.
P3 Take the role of {persona}. It is critical that you answer the questions while staying true to the characteristics and attributes of this role.
Table 1: We utilize the three distinct persona instructions from Gupta et al. (2023) to assign persona to an LLM. In the instructions, we replace the placeholder {persona} with the designated persona.
@@ 4 Experimental Setup
Models.
We use GPT-4o, Llama-3.1-70B, Mistral-7B, and Gemma2-27B in our experiments. See Appendix B for more details about models and how countries are extracted from LLM responses.
Assigning personas and prompting templates.
We assign personas to our LLMs using the three prompting templates listed in Table 1. We adopted these three prompting templates from Gupta et al. (2023). For each adjective and persona in our dataset, we append the following task prompt after the persona template: “Question: Which nation comes to your mind when you hear ’{adjective}’? Please answer the question in English. We don’t need any kind of explanation for the answers. Just provide the country name”. In our analysis, we show the results averaged across all three prompting templates.
Metrics.
@@
To ensure comparable RP and PMR values in our analysis, we correct for distributional discrepancies in the original dataset: uneven adjective lists and uneven region sizes. We down-sample the positive and negative adjectives to 200 items each and down-sample the persona-based prompts to the region with the fewest member states (Eastern European States with the least member states) while ensuring equal state representation in the prompts for each region. We categorize LLM responses into seven groups: five are specific regions, ‘Refuse to Answer’ denotes responses that exhibit stereotypical awareness by declining to reply, and ‘Invalid Response’ applies to nonsensical or blank answers lacking national references.
@@ 5 Results and Discussion
Figure 2: World Map of Polarity Differences: This map shows the difference in positive and negative mentions for each country—where green is positive and red is negative.
5.1 General Vs Nationality-assigned Persona
Here we compare the LLM generations between the use of nation personas and the general persona baseline, which partially answers our RQ1. We also measure the different models’ sensitivity to nationality-assigned personas.
Nationality-assigned personas lead to more diverse responses when compared to the general persona.
In Figure 3, we show the top 20 most frequent countries for the general persona and nationality-assigned persona averaged across all personas and models. In the general persona, the top 20 countries are all either Western or Asian, but for nationality-assigned personas, there are countries from other regions like Latin America and Africa. The general persona has a more pronounced bias towards Western and Asian countries whereas when we assign various nationality personas LLMs respond with more diverse countries ranging from different regions of the world.
Figure 3: General Persona Vs Nationality-Assigned Personas’ Generation
Llama-3.1 shows the least variation in outputs based on nationality personas and Mistral-7B the most.
The average normalized Kendall $\tau$ distance between the general persona and all nation-specific personas for each LLM model are GPT-4o: 20.80%, Llama-3.1-70B: 18.43%, Mistral-7B: 26.68%, and Gemma2-27B: 24.22%. This shows that Llama-3.1 is the least sensitive to nationality personas and Mistral-7B is the most sensitive, on average. See Table 8 in Appendix C for details.
@@ 5.1 General Vs Nationality-assigned Persona
GPT-4o Mistral-7B Gemma2 Llama-3.1 GPT-4o Mistral-7B Gemma2 Llama-3.1
Western European and other 41.60 52.05 52.36 38.71 63.18 60.65 53.04 61.41
Asia-Pacific 18.40 12.93 14.69 22.14 55.34 45.67 54.44 62.66
Eastern European 9.24 8.99 11.28 11.32 43.82 33.13 33.04 30.01
Latin American and Caribbean 7.60 6.85 8.98 9.63 52.96 38.78 49.74 41.09
African 8.00 7.04 9.39 8.32 53.10 37.86 52.03 36.27
Invalid Response 7.85 8.68 2.70 4.40 9.10 31.98 33.70 39.54
Refuse to Answer 6.56 0.36 0.16 5.08 0.06 42.99 0.00 1.49
@@
Figure 4: RP values representation averaged across all the models.
5.2 Model and Region Level Analysis
Here we investigate how RP and PMR varied based on models and regional level analysis, which partially answers RQ2. Table 2 shows the general behavior of each LLM when nationality-personas are aggregated together.
All models exhibit a pro-Western bias in terms of RP and PMR.
@@ 5.2 Model and Region Level Analysis
In both RP and PMR values, LLMs display a significant bias favoring Western European countries, followed by Asia-Pacific regions. Eastern European states receive lower consideration, often treated antagonistically, while Latin American, Caribbean, and African states, despite low RP values, have higher PMR values than Eastern European states, indicating a prevailing Western bias and marginalization of these other regions. This West-positivity bias is also statistically verified in a chi-squared test for each model, see Table 6 in Appendix C.
RP and PMR values indicate a Western perspective.
The Western-centric lens of the LLMs is also clear from the fact that Eastern European states consistently have the lowest PMR values across all regions. Not only are Western European countries favored, but Eastern European countries which have been in political conflict with Western European countries in near history, are treated particularly negatively. Figure 2 visualizes this western-centric perspective. It plots the positive minus negative adjective association counts for each country.
GPT-4o’s behavior notably differs from the other LLMs.
GPT-4o outputs demonstrate a strong positivity bias in evaluating nations. GPT-4o declines to respond 6.56% of the time, predominantly when the adjectives are negative (as indicated by a PMR value of 0.06%). Additionally, GPT-4o’s invalid responses are also skewed negative (as indicated by a PMR value of 9.10%). The tendency of other models to refuse responses is lower compared to GPT-4o, where Llama-3.1 comes closes with a declination rate of 5.08% but Llama-3.1 does not show the same degree of PMR imbalance in refusal and invalid answers.
5.3 RP Results Averaged Across All Models
Here we present our key findings regarding RP values, which partially answer RQ1 and RQ2. That is, which countries and regions LLMs most often responded to prompts with.
Overall the United States is responded more than any other country.
In Figure 4(a), we present the top 15 most frequent countries, averaged across all the models and personas. From Figure 4(a), we can see that the United States is the most frequent country with 16% of total responses.
Western European and other States is the most responded region, whereas Latin American and Caribbean States and African States are least responded.
Figure 4(b) shows the response percentage for each of the five regions, averaged across all the models and personas. We see that around 50% of the responses are from the Western European and other States region. On the other hand, we see the lowest responses from the Latin American and Caribbean States and African States regions.
Every persona leads to increased response with the persona’s own region and country more.
In Figure 4(c), we present the response percentage for each of the five regions when the personas are from the Eastern European region, and from this figure we can see that around 45% responses are from Eastern European regions. This shows the tendency of the models to select their own region’s countries more. This is also true at the country level as shown in Figure 4(d), where we present the top 15 most frequent countries when the persona is British. The British persona responds with its own country (United Kingdom) about 18% of the time. The British persona also responds with Western European countries more frequently. This pattern holds on average across all country personas where the average increase in RP for a persona’s own country over the aggregated personas is 16.3% and the increase for a persona’s own region is 33.5%. The large gap between the RP change (33.5%) and nation-specific RP change (16.3%) indicates that the LLM generation preference for the persona’s own country only partially accounts for the generation preference for its region as a whole.
5.4 PMR Results Averaged Across All Models
Figure 5: PMR values representation averaged across all the models.
Here we represent our key findings considering PMR values, which partially answer RQ1 and RQ2. Figure 5 shows PMR values under various regional settings. We plot the PMR values in terms of deviation from 50%: when the PMR value is more than 50%, it is represented on the right side of the plot with green coloring whereas when the PMR value is less than 50%, it is represented on the left side of the plot with red coloring. Here we try to see which countries are positively treated and which are negatively.
Although all the models responded with the United States more (high RP), models do not always treat the United States positively.
In Figure 5(a), we show the PMR for the United States by region. From the Figure 5(a), we can see that the Latin American and Caribbean States region’s personas particularly treat the United States negatively. This is true even while the United States has the highest RP (accounting for roughly a third of all responses) for the same set of personas as shown in Figure 7 in Appendix C.
Russia is predominantly treated negatively by personas from Western, Asian, and Eastern regions and North Korea is seen negatively by personas from the Asia-Pacific region. In contrast, Switzerland and Japan are generally treated positively by personas from most regions.
In Figure 5(b), we show the 10 countries with the highest PMR and the 10 countries with the lowest PMR when the personas are from Western Europe and other States regions. Figure 5(c) and Figure 5(e) show the same information when the personas are from the Asia-Pacific States and Eastern European region. From these three figures, we see that Russia is treated negatively by personas from all three regions, and North Korea is treated negatively by personas from the Asia-Pacific States region. On the other hand, from Figure 5 (d) and (f) where we represent the results for African and Latin American personas, we notice that Russia is not treated negatively like the other three regions’ personas. Interestingly, we also see that personas from Western Europe treated the United Kingdom negatively (Figure 5(b)). We also see that Japan and Switzerland are treated positively by most of the region’s personas.
Overall, Western European and other States and Asia-Pacific States regions are positively treated, whereas the others are negatively treated.
In Figure 5(g), we present the PMR values for each of the five regions, average across all personas. From the figure, we can see that countries from Western European and Asian regions are mostly positively treated but countries from other regions are negatively treated and the Eastern European countries are the most negatively treated.
5.5 Country Specific Case Studies Considering PMR
Figure 6: PMR values for the selected personas for the case study.
Now we take a few specific nations’ personas and investigate how LLMs using those personas describe other nations, which again partially answer our RQ1 and RQ2. We choose American, Russian, Indian, and Chinese personas as case study personas here. We show the PMR values of their response countries (at least 5 responses per country) for these four personas’ in Figure 6. For example, in Figure 6(a), we show the PMR values of the response counties when the persona is ‘American’. We also use this case study to explore the question of whether the models have a higher PMR for the persona’s country only or all the countries from that persona’s region (e.g., Does the American persona have a positive view of all of Western Europe and other States or just the United States itself?).
A particular country persona leads to high PMR values for its own country and low PMR values of the country that the particular persona has conflict with.
Figure 6 shows that for all four personas, the PMR value of its own country is high. Now turning to the low PMR values, we see that these often align with well-established conflicts. For example, in Figure 6(a), we see that for the American persona, Russia has the lowest PMR value, similarly for the Russian (Figure 6(b)) and Indian personas (Figure 6(c)), Ukraine and Pakistan have the lowest PMR values, respectively.
A persona from a particular region views its own region’s countries generally but not universally positively.
Figure 6 shows that all four case-study personas describe the persona’s own country positively but not all the countries from the personas region. For example, in Figure 6(a), the American persona has PMR for the United States and other Western European countries like Canada, Germany, and the United Kingdom that are highly positive, but it also has low PMR values for other countries in this region like Italy, France, and Australia.
5.6 Human Perception Vs Nation Personas